OpenStudy (anonymous):

f(x) = 9x^3 − 27x − 19 x = I got 1, but the homework tells me it is wrong, PLEASE help me out!

OpenStudy (anonymous):

\[27x ^{3} - 27 = 0 and then 27x ^{2} = 27 x ^{2} = 1 x = 1\]

OpenStudy (anonymous):

27x^2 - 27 = 0

OpenStudy (anonymous):

27x^2 = 27

OpenStudy (anonymous):

x^2 = 1

OpenStudy (anonymous):

so therefore x = 1????

OpenStudy (anonymous):

could it be (-1,1)

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

on the other problems they drop the 19 and just use chain rule and solve for x

OpenStudy (anonymous):

power*

OpenStudy (anonymous):

no, -3,3 isnt right either, im thinking about telling my teacher that she is wrong. im serious ive been in a library for three hours doing homework!

OpenStudy (anonymous):

what do you get for this one f(x) = 4x^3 + 4x^−3

OpenStudy (anonymous):

you're finding critical points